Question 1 The engineering team at Manuels Manufacturing, Inc., is planning to purchase an enterprise resource planning (ERP) system. The software and installation from Vendor A costs $430,000 initially and is expected to increase revenue $110,000 per year every year. The software and installation from Vendor B costs $285,000 and is expected to increase revenue $115,000 per year. Manuels uses a 4- year planning horizon and a 7.5 % per year MARR. Click here to access the TVM Factor Table Calculator What is the present worth of each investment? Vendor A: $ Vendor B: $ Carry all interim calculations to 5 decimal places and then round your final answer to the nearest dollar. The tolerance is ±20. What is the decision rule for determining the preferred investment based on present worth ranking? Which ERP system should Manuel purchase?
Answer #1

P/A(r%, N) = [1 - (1 + r)-N] / r

P/A(7.5%, 4) = [1 - (1.075)-4] / 0.075 = (1 - 0.7488) / 0.075 = 0.2512 / 0.075 = 3.3493

Therefore,

(1)

PW, Vendor A ($) = - 430,000 + 110,000 x P/A(7.5%, 4) = - 430,000 + 110,000 x 3.3493 = - 430,000 + 368,423

= - 61,577

PW, Vendor B ($) = - 285,000 + 115,000 x P/A(7.5%, 4) = - 285,000 + 115,000 x 3.3493 = - 285,000 + 385,170

= 100,170

(2) Decision rule is:

Accept project if PW > 0

Reject project if PW < 0

(3) Manual should purchase from Vendor B (Since PW > 0).